Please answer the following question. Please show all your working/solutions.

In dealing with macroeconomic data, it is often informative to express GDP- per-capita in logs. To see the convenience, consider a variable Xt over time. The growth rate of a variable Xt from period t − 1 to period t is given by (XtXt−1)/Xt−1. If we let (Xt) denote the growth rate of the variable Xt from t − 1 to t, we can say that ∆(Xt) is well approximated by log(Xt/Xt−1). Let’s see why.

(a) Show that for small values of z, we must have log(1 + z) ≈ z. (Hint: Use Taylor expansion!). Use this result in the following parts.

(b) Show that log(Xt/Xt−1)(Xt − Xt−1)/Xt−1 = ∆(Xt). (Hint: add and subtract 1 to Xt/Xt−1 inside the log function.) For the rest of the question, assume ∆(Xt) = log(Xt/Xt−1).

(c) Show ∆(XtYt) = ∆(Xt) + ∆(Yt)

(d) Show ∆(Xt/Yt) = ∆(Xt) − ∆(Yt)

(e) Show ∆(Xt2) = 2∆(Xt)
Answer #1

Ans A)

Log(1+z) with Taylor expansion can be written as follows

Log(1+z)=z+z^2/2+z^3/3+...

As z tends to 0

Higher order of z tends to 0

Hence Log(1+z)=z (approximately)

Ans B)

Log(1+((Xt/Xt-1)-1))=(Xt/Xt-1)-1 (approx.) As per proof above

Now

Log(Xt/Xt-1)=LogXt-LogXt-1=(Log(1+Xt-1)-Log(1+Xt-1-1))=

Xt-1+(Xt-1)^2/2+(Xt-1)^3/3+...-(Xt-1-1)-(Xt-1-1)^2/2-...=Xt-Xt-1

If we ignore higher order terms of Xt and Xt-1

Log(Xt/Xt-1)=∆Xt

Ans C)

∆(Xt)+∆(Yt)=log(Xt/Xt-1)+log(Yt/Yt-1)=log(XtYt/(Xt-1)(Yt-1))=∆(XtYt)

Ans D)

∆(Xt)-∆(Yt)=log(Xt/Xt-1)-log(Yt/Yt-1)=log(Xt/(Xt-1)/Yt/(Yt-1))=log((Xt/Yt)/(Xt-1/Yt-1))=∆(Xt/Yt)

Ans E)

∆(Xt^2)=log(Xt^2/Xt-1^2)=Log((Xt/Xt-1)^2)=2Log((Xt/Xt-1))=2∆Xt